### v1

@@ 2.3 The strength of the self-interactions
A remaining question is how close a given cross-section would be to the collisionless or gaseous regime. This means we need to understand how much the trajectory of the DM particles is impacted by gravity compared to the scattering. A common measure for this in the context of SIDM halos is the Knudsen number (e.g. Koda & Shapiro, 2011). It is given by
$\mathrm{Kn}=\sqrt{\frac{4\pi\mathrm{G}}{\rho\sigma_{\mathrm{DM}}^{2}}}\left(\frac{\sigma_{\mathrm{V}}}{m}\right)^{-1}\,.$ (10)
This is comparing the ratio of the mean free path to the Jeans length. The latter is a measure for how susceptible the DM medium is to gravitational collapse. We note that the Knudsen number, Kn, is independent of the mass of the perturber, i.e., the BH mass. However, the strength of the dynamical friction depends on it. For Eq. 10 we used the viscosity cross-section per mass, $\sigma_{\mathrm{V}}/m$, for defining the Knudsen number. It is given by
@@
Here, we follow the normalisation of Yang & Yu (2022) implying that viscosity and total cross-section have equal values for isotropic scattering ($\left.\sigma_{\mathrm{V}}\right|_{\rm iso}=\left.\sigma_{\mathrm{tot}}\right|_{\rm iso}$). The viscosity cross-section has the advantage that it often provides a good matching between different angular dependencies (e.g., Yang & Yu, 2022; Colquhoun et al., 2021) and is thus a good characteristic for the strength of the scattering. If the Knudsen number is large, i.e. $\mathrm{Kn}\gg 1$, a system is considered to be in the long-mean-free-path regime, i.e., gravity dominates over the scattering. In contrast, if $\mathrm{Kn}\ll 1$, the system would be in the short-mean-free-path regime and scattering dominates over gravity. So we expect that the system is close to the collisionless case when the Knudsen number is larger, and for small numbers falls into the gaseous regime.
For the DM spike, the Knudsen number becomes smaller when getting closer to the primary BH as density and velocity dispersion increase. This implies that the behaviour of the dynamical friction could transition from the collisionless regime to the gaseous one while the secondary BH is inspiralling. However, it depends on the strength of the DM self-interactions.
Last, we note that also in the case of a primary BH hosting a DM spike, the heat conduction due to self-interactions being in the long-mean-free-path or short-mean-free-path regime is independent of the mass of the primary BH. In other words, the Knudsen number does not depend on the mass of the primary BH. We note that this is in contradiction to the assumption by Shapiro & Paschalidis (2014); Shapiro (2018). In consequence, a DM spike given a cross-section within the typically studied range of $\approx$ 0.1–100 cm2g-1 (e.g. Tulin & Yu, 2018; Adhikari et al., 2022) would be to a large extent rather in the short-mean-free-path regime than in the long-mean-free-path regime. This is in line with the work by Pollack et al. (2015), where they studied a DM model with a fraction of the matter being ultra-strongly self-interacting. Their heat conductivity does only depend on the density of the ultra-strongly interacting matter component.
@@ 3.1 Simulation Code
We use the cosmological $N$-body code OpenGadget3 (see Groth et al., 2023, and the references therein). The SIDM module of the code has been described by Fischer et al. (2021a, b, 2022, 2024b). Note, it can simulate both large and small-angle scatterings. To the first one, we also refer to as rare scattering since a single scattering event on average leads to a larger energy and momentum transfer between the particles compared to small-angle scattering, which must be much more frequent to have a similar effect. For the first one, the angular dependence is modelled explicitly by drawing random angles with respect to the differential cross-section, for the second one an effective description is used based on a drag-like behaviour (see also Kahlhoefer et al., 2014). The interactions between the numerical particles are modelled on a pair-wise basis. We compute the scattering probability for the rarely self-interacting dark matter (rSIDM) and the drag force for the frequently self-interacting dark matter (fSIDM) using a spline kernel (Monaghan & Lattanzio, 1985). Its size is determined with the next neighbour approach using $N_{\mathrm{ngb}}=64$ neighbours. This choice is known to provide accurate results (Fischer et al., 2021a). We employ a separate time step criterion for SIDM to keep the interaction probability per interaction sufficiently low. The SIDM module is designed to explicitly conserve linear momentum and energy, even when executed in parallel.
To model the gaseous regime when self-interactions are very strong, we use smoothed-particle hydrodynamics (SPH, Gingold & Monaghan, 1977). Compared to gadget-2 (Springel, 2005), an updated SPH formulation (Beck et al., 2016) with an updated viscosity scheme (Dolag et al., 2005; Beck et al., 2016) is employed. We choose a Wendland $C^{6}$ kernel (Dehnen & Aly, 2012) with its size being determined by the, $N_{\mathrm{ngb}}=230$, next neighbours.
@@ 3.2 Simulation Setup
To test the dynamical friction, we simulate a BH travelling through a constant density consisting of DM. The mass density is set to $\rho=4.815\times 10^{-4}\,\mathrm{M_{\odot}}\,{\mu\mathrm{pc}}^{-3}$. The velocities of the DM particles follow a Maxwell-Boltzmann distribution, and their one-dimensional velocity dispersion is chosen to be $\sigma_{\mathrm{1D}}=625.5\,\mathrm{km}\,\mathrm{s}^{-1}$. The BH is represented by a single particle with a mass of $m_{\mathrm{BH}}=7\times 10^{3}\,\mathrm{M_{\odot}}$. Initially, it is moving with a velocity of $v_{\mathrm{ini}}=1037\,\mathrm{km}\,\mathrm{s}^{-1}$. These values are selected because they lead to a strong dynamical friction force, and we expect to find a significant impact in our simulations. Furthermore, we chose a cubic simulation volume with a side length of $l_{\mathrm{box}}=829.5\,\mu\mathrm{pc}$ for most of our simulations but also use a box length of $l_{\mathrm{box}}=663.6\,\mu\mathrm{pc}$. We vary the box length to study its impact on the dynamical friction force. In addition, periodic boundary conditions are employed. The DM is resolved by $N_{\mathrm{DM}}=5\times 10^{6}$ particles, which leads to a numerical particle mass of $m_{\mathrm{DM}}=5.5\times 10^{-2}\,\mathrm{M_{\odot}}$. For testing the accuracy of our simulations, we also perform runs with $N_{\mathrm{DM}}=1\times 10^{6}$ DM particles, i.e., a five times higher particle mass. We set the Plummer-equivalent gravitational softening length to $\epsilon=24\,\mu\mathrm{pc}$. The gravitational forces are computed between all particles, i.e., we also take the self-gravity of the perturbed medium into account. In principle, the self-gravity can be relevant for the resulting dynamical friction (see Just & Kegel, 1990). However, we do not separately study its contribution.
Note, that the $N$-body representation of the constant density is subject to fluctuation due to the sampling noise. The fluctuations can collapse under gravitational influence. We have checked that, for our choice of the box length and resolution, this does not occur within the time we simulate.
We run simulations with a very anisotropic cross-section by employing the numerical scheme for frequent self-interactions developed by Fischer et al. (2021a). Here the strength of the self-interactions is specified in terms of the momentum transfer cross-section,
@@
We simulate velocity-independent elastic scattering with cross-sections of $\sigma_{\mathrm{V}}/m\in\{3\times 10^{-9},3\times 10^{-8},3\times 10^{-7},9\times 10^{-7},3\times 10^{-6}\}\,(\mathrm{cm}^{2}\,\mathrm{g}^{-1})$. [^1] To make it easier to compare cross-sections of different angular dependencies, we do not specify the cross-section in terms of the momentum transfer cross-section but use the viscosity cross-section (see Eq. 11). In addition, we simulate velocity-independent isotropic cross-sections with a total cross-section of $\sigma_{\mathrm{V}}/m\in\{2\times 10^{-8},3\times 10^{-8},2\times 10^{-7},3\times 10^{-7}\}\,(\mathrm{cm}^{2}\,\mathrm{g}^{-1})$ to investigate the role of the angular dependence. We summarise all the cross-sections used in Tab. 1.
In addition, we also investigate velocity-dependent self-interactions. Here, we use the implementation by Fischer et al. (2024b). We parameterise the velocity dependence of the viscosity cross-section as
@@
For the velocity parameter, $w$, we choose $w\in\{300,461.1,683.2,1483.5\}\,(\mathrm{km}\,\mathrm{s}^{-1})$. Note, the mean relative velocity of the DM particles in the ICs is $\langle v\rangle=1411.6\,\mathrm{km}\,\mathrm{s}^{-1}$. For all simulations with a velocity-dependent cross-section, we choose the value of $\sigma_{0}$ to result in an effective cross-section (see Yang & Yu, 2022) of $\sigma_{\mathrm{eff}}/m=3\times 10^{-8}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. A summary of all cross-sections is given in Tab. 1.
Note, that all the cross-sections mentioned above are many orders of magnitude smaller than the tightest constraints on SIDM in the current literature (e.g. Harvey et al., 2019; Andrade et al., 2021; Sagunski et al., 2021; Eckert, D. et al., 2022; Gopika & Desai, 2023). Cross-sections as large as the ones typically discussed in the astrophysical context imply a mean free path substantially shorter than the kernel size, and may no longer be faithfully simulated with the standard SIDM schemes (see our Appendix A as well as Fischer et al., 2024a). In addition, the simulations become more costly with increasing cross-section as the required time-step decreases. To model the regime of such a large cross-section, we use SPH. For the ICs, we set the initial velocity of the SPH particles to zero and choose their internal energy to correspond to the same velocity dispersion as in the SIDM setup.
@@
Table 1: The table gives the different cross-sections we use for the simulations of our test setup. The first column specifies the type of scattering, where frequent refers to a very anisotropic cross-section and rare to an isotropic one. The second row gives the viscosity cross-section (Eq. 11). In the case of the velocity-dependent cross-sections, the second row corresponds to $\sigma_{0}/m$ of Eq. (13). The last column gives the parameter $w$ of Eq. (13) for the velocity-dependent cross-sections.
@@ 4 Test Problems
In this section, we present the results from our simulations following the setup described in section 3. We begin with the collisionless case and investigate the impact of variations in the numerical setup. Subsequently, we study the role of the strength of the self-interactions and its impact on the density wake. Furthermore, we investigate the matching of different angular and velocity dependencies onto each other in the case of dynamical friction.
@@ 4.1 The collisionless case
Figure 3: The velocity and position of the BH in our test setup are shown as a function of time. All simulations are for collisionless DM. The solid lines are from the setup with a box length of $l=829.5\mu\mathrm{pc}$. In contrast, the dashed lines correspond to a setup with a smaller box size of $l=663.6\mu\mathrm{pc}$. The high-resolution results correspond to $N_{\mathrm{DM}}=5\times 10^{6}$ DM particles, and the low-resolution ones to $N_{\mathrm{DM}}=10^{6}$. Note that in the smaller box setup, the low resolution implies a higher mass resolution than in the large box setup.
@@
In our test setup as described in Section 3, the BH is travelling through an initially constant density. The BH perturbs the DM and creates a density wake, leading effectively to a force decelerating the BH. This dynamical friction force depends on the velocity of the BH relative to the velocity dispersion of the perturbed DM density (see Fig. 2). In our setup, we first expect the dynamical friction to become stronger while the BH is slowing down. After the BH velocity has decreased enough, it enters the phase where the dynamical friction becomes weaker with decreasing velocity.
@@
In Fig. 3, we show the position and velocity of the BH in our test setup for collisionless matter. As expected, the velocity of the BH decreases over the course of the simulation. The results for two different box sizes are shown. We find that the dynamical friction is reduced for a smaller box size. This is not surprising, as $r_{\mathrm{max}}$ in the Coulomb logarithm of the Chandrasekhar formula, Eq. (7), depends for our setup effectively on the box size because the box size limits the maximal possible impact parameter. The results agree qualitatively with the Chandrasekhar formula, after an initial phase during which the density wake forms, i.e., Eq. (7) provides a good description when the Coulomb logarithm is taken as a free parameter. For the large box, we show the results using two different resolutions. In both cases, the trajectory of the BH is very small, such that we can safely assume that the resolution for our study is large enough.

### v2

@@ 2.3 The strength of the self-interactions
A remaining question is how close a given cross-section would be to the collisionless or gaseous regime. This means we need to understand how much the trajectory of the DM particles is impacted by gravity compared to the scattering. A common measure for this in the context of SIDM haloes is the Knudsen number (e.g. Koda & Shapiro, 2011). It is given by
$\mathrm{Kn}=\sqrt{\frac{4\pi\mathrm{G}}{\rho\sigma_{\mathrm{DM}}^{2}}}\left(\frac{\sigma_{\mathrm{eff}}}{m}\right)^{-1}\,.$ (10)
This is comparing the ratio of the mean free path to the Jeans length. The latter is a measure for how susceptible the DM medium is to gravitational collapse. We note that the Knudsen number, Kn, is independent of the mass of the perturber, i.e. the BH mass. However, the strength of the dynamical friction depends on it. For Eq. 10 we used the effective cross-section per mass, $\sigma_{\mathrm{eff}}/m$, for defining the Knudsen number. The effective cross-section (Yang & Yu, 2022) allows one to map cross-sections with different velocity dependencies onto each other and is given by
Where an average integrating over the distribution of relative velocities, $v$, and weighting the cross-section with $v^{5}$ is computed assuming that the relative velocities follow a Maxwell–Boltzmann distribution. Furthermore, $\sigma_{\mathrm{V}}$ is the viscosity cross-section,
@@
Here, we follow the normalisation of Yang & Yu (2022) implying that viscosity and total cross-section have equal values for isotropic scattering ($\left.\sigma_{\mathrm{V}}\right|_{\rm iso}=\left.\sigma_{\mathrm{tot}}\right|_{\rm iso}$). The viscosity cross-section has the advantage that it often provides a good matching between different angular dependencies (e.g. Yang & Yu, 2022; Colquhoun et al., 2021) and is thus a good characteristic for the strength of the scattering. If the Knudsen number is large, i.e. $\mathrm{Kn}\gg 1$, a system is considered to be in the long-mean-free-path regime, i.e. gravity dominates over the scattering. In contrast, if $\mathrm{Kn}\ll 1$, the system would be in the short-mean-free-path regime and scattering dominates over gravity. So we expect that the system is close to the collisionless case when the Knudsen number is larger, and for small numbers falls into the gaseous regime.
For the DM spike, the Knudsen number becomes smaller when getting closer to the primary BH as density and velocity dispersion increase. The Knudsen number scales as $\mathrm{Kn}\propto r^{(1+\alpha_{\mathrm{sp}})/2}$ when assuming a density spike profile as given by Eq. (1) and the gravitational potential being dominated by the BH. This implies that the behaviour of the dynamical friction could transition from the collisionless regime to the gaseous one while the secondary BH is inspiralling. However, it depends on the strength of the DM self-interactions.
Last, we note that also in the case of a primary BH hosting a DM spike, the heat conduction due to self-interactions being in the long-mean-free-path or short-mean-free-path regime is independent of the mass of the primary BH if Eq. 10 holds. In other words, the Knudsen number does not depend on the mass of the primary BH. We note that this is in contradiction to the assumption by Shapiro & Paschalidis (2014) and Shapiro (2018). Instead of using the Jeans length for the Knudsen number, they approximated the gravitational scale height as the distance, $r$, to the primary BH. In fact, their Knudsen number is given as the mean free path divided by $r$. The influence of the BH changing the gravitational scale height can be motivated by studies of star clusters (e.g. Bettwieser & Spurzem, 1986). It is unclear if this is in general the same for the SIDM case. If instead Eq. 10 holds, a DM spike given a cross-section within the typically studied range of $\approx$ 0.1–100 cm2g-1 (e.g. Tulin & Yu, 2018; Adhikari et al., 2022) could be to a large extent rather in the short-mean-free-path regime than in the long-mean-free-path regime. This would be in line with the work by Pollack et al. (2015), where they studied a DM model with a fraction of the matter being ultra-strongly self-interacting. Their heat conductivity does only depend on the density of the ultra-strongly interacting matter component.
@@ 3.1 Simulation code
We used the cosmological $N$-body code OpenGadget3 (see Groth et al., 2023, and the references therein). The SIDM module of the code has been described by Fischer et al. (2021a, b, 2022, 2024b). We employed it to model DM interactions beyond gravity. The code can simulate both large and small-angle scatterings. To the first one, we also refer to as rare scattering since a single scattering event on average leads to a larger energy and momentum transfer between the particles compared to small-angle scattering, which must be much more frequent to have a similar effect. For the first one, the angular dependence is modelled explicitly by drawing random angles with respect to the differential cross-section, for the second one an effective description is used based on a drag-like behaviour (see also Kahlhoefer et al., 2014). The interactions between the numerical particles are modelled on a pair-wise basis. We have computed the scattering probability for the rarely self-interacting dark matter (rSIDM) and the drag force for the frequently self-interacting dark matter (fSIDM) using a spline kernel (Monaghan & Lattanzio, 1985). Its size is determined with the next neighbour approach using $N_{\mathrm{ngb}}=64$ neighbours. This choice is known to provide accurate results (Fischer et al., 2021a). We employed a separate time step criterion for SIDM to keep the interaction probability per interaction sufficiently low. The SIDM module is designed to explicitly conserve linear momentum and energy, even when executed in parallel.
To model the gaseous regime when self-interactions are very strong, we used smoothed-particle hydrodynamics (SPH, Gingold & Monaghan, 1977). Compared to gadget-2 (Springel, 2005), an updated SPH formulation (Beck et al., 2016) with an updated viscosity scheme (Dolag et al., 2005; Beck et al., 2016) was employed. We chose a Wendland $C^{6}$ kernel (Dehnen & Aly, 2012) with its size being determined by the, $N_{\mathrm{ngb}}=230$, next neighbours.
@@ 3.2 Simulation setup
To test the dynamical friction, we simulated a BH travelling through a constant density consisting of DM. The mass density is set to $\rho=4.815\times 10^{-4}\,\mathrm{M_{\odot}}\,{\mu\mathrm{pc}}^{-3}$. The velocities of the DM particles follow a Maxwell-Boltzmann distribution, and their one-dimensional velocity dispersion is chosen to be $\sigma_{\mathrm{1D}}=625.5\,\mathrm{km}\,\mathrm{s}^{-1}$. The BH is represented by a single particle with a mass of $m_{\mathrm{BH}}=7\times 10^{3}\,\mathrm{M_{\odot}}$. Initially, it is moving with a velocity of $v_{\mathrm{ini}}=1037\,\mathrm{km}\,\mathrm{s}^{-1}$. These values are selected because they lead to a strong dynamical friction force, and we expect to find a significant impact in our simulations. Furthermore, we chose a cubic simulation volume with a side length of $l_{\mathrm{box}}=829.5\,\mu\mathrm{pc}$ for most of our simulations but also used a box length of $l_{\mathrm{box}}=663.6\,\mu\mathrm{pc}$. We varied the box length to study its impact on the dynamical friction force. In addition, periodic boundary conditions were employed. We note that this also implies a periodic computation of the gravitational forces. Otherwise, the distribution of DM would easily collapse. The DM is resolved by $N_{\mathrm{DM}}=5\times 10^{6}$ particles, which leads to a numerical particle mass of $m_{\mathrm{DM}}=5.5\times 10^{-2}\,\mathrm{M_{\odot}}$. For testing the accuracy of our simulations, we also performed runs with $N_{\mathrm{DM}}=1\times 10^{6}$ DM particles, i.e. a five times higher particle mass. We set the Plummer-equivalent gravitational softening length to $\epsilon=24\,\mu\mathrm{pc}$. The gravitational forces are computed between all particles, i.e. we also take the self-gravity of the perturbed medium into account. In principle, the self-gravity can be relevant for the resulting dynamical friction (see Just & Kegel, 1990). However, we do not separately study its contribution.
We note that the $N$-body representation of the constant density is subject to fluctuation due to the sampling noise. The fluctuations can collapse under gravitational influence. We have checked that, for our choice of the box length and resolution, this does not occur within the time we simulate.
We ran simulations with a very anisotropic cross-section by employing the numerical scheme for frequent self-interactions developed by Fischer et al. (2021a). Here the strength of the self-interactions is specified in terms of the momentum transfer cross-section,
@@
We simulated velocity-independent elastic scattering with cross-sections of $\sigma_{\mathrm{V}}/m\in\{3\times 10^{-9},3\times 10^{-8},3\times 10^{-7},9\times 10^{-7},3\times 10^{-6}\}\,(\mathrm{cm}^{2}\,\mathrm{g}^{-1})$. [^1] To make it easier to compare cross-sections of different angular dependencies, we do not specify the cross-section in terms of the momentum transfer cross-section but use the viscosity cross-section (see Eq. 12). In addition, we simulated velocity-independent isotropic cross-sections with a total cross-section of $\sigma_{\mathrm{V}}/m\in\{2\times 10^{-8},3\times 10^{-8},2\times 10^{-7},3\times 10^{-7}\}\,(\mathrm{cm}^{2}\,\mathrm{g}^{-1})$ to investigate the role of the angular dependence. We summarise all the cross-sections used in Table 1.
In addition, we also investigated velocity-dependent self-interactions. Here, we used the implementation by Fischer et al. (2024b). We parameterise the velocity dependence of the viscosity cross-section as
@@
For the velocity parameter, $w$, we choose $w\in\{300,461.1,683.2,1483.5\}\,(\mathrm{km}\,\mathrm{s}^{-1})$. We note that the mean relative velocity of the DM particles in the ICs is $\langle v\rangle=1411.6\,\mathrm{km}\,\mathrm{s}^{-1}$. For all simulations with a velocity-dependent cross-section, we chose the value of $\sigma_{0}$ to result in an effective cross-section (see Yang & Yu, 2022) of $\sigma_{\mathrm{eff}}/m=3\times 10^{-8}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. A summary of all cross-sections is given in Table 1.
We note that all the cross-sections mentioned above are many orders of magnitude smaller than the tightest constraints on SIDM in the current literature (e.g. Harvey et al., 2019; Andrade et al., 2021; Sagunski et al., 2021; Eckert, D. et al., 2022; Gopika & Desai, 2023). Cross-sections as large as the ones typically discussed in the astrophysical context imply a mean free path substantially shorter than the kernel size, and may no longer be faithfully simulated with the standard SIDM schemes (see our Appendix A as well as Fischer et al., 2024a). In addition, the simulations become more costly with increasing cross-section as the required time-step decreases. To model the regime of such a large cross-section, we use SPH. For the ICs, we set the initial velocity of the SPH particles to zero and chose their internal energy to correspond to the same velocity dispersion as in the SIDM setup.
Table 1: Simulated cross-sections.
@@
 [^2] The table gives the different cross-sections we used for the simulations of our test setup. The first column specifies the type of scattering, where frequent refers to a very anisotropic cross-section and rare to an isotropic one. The second column gives the viscosity cross-section (Eq. 12). In the case of the velocity-dependent cross-sections, the second column corresponds to $\sigma_{0}/m$ of Eq. (14). The last column gives the parameter $w$ of Eq. (14) for the velocity-dependent cross-sections.
@@ 4 Test problems
In this section, we present the results from our simulations following the setup described in Sect. 3. We begin with the collisionless case and investigate the impact of variations in the numerical setup. Subsequently, we study the role of the strength of the self-interactions and its impact on the density wake. Furthermore, we investigate the matching of different angular and velocity dependencies onto each other in the case of dynamical friction.
@@ 4.1 The collisionless case
Figure 3: Velocity and position of the BH in our test setup as a function of time. All simulations are for collisionless DM. The solid lines are from the setup with a box length of $l=829.5\mu\mathrm{pc}$. In contrast, the dashed lines correspond to a setup with a smaller box size of $l=663.6\mu\mathrm{pc}$. The high-resolution results correspond to $N_{\mathrm{DM}}=5\times 10^{6}$ DM particles, and the low-resolution ones to $N_{\mathrm{DM}}=10^{6}$. We note that in the smaller box setup, the low resolution implies a higher mass resolution than in the large box setup.
@@
In our test setup as described in Sect. 3, the BH is travelling through an initially constant density. The BH perturbs the DM and creates a density wake, leading effectively to a force decelerating the BH. This dynamical friction force depends on the velocity of the BH relative to the velocity dispersion of the perturbed DM density (see Fig. 2). In our setup, we first expect the dynamical friction to become stronger while the BH is slowing down. After the BH velocity has decreased enough, it enters the phase where the dynamical friction becomes weaker with decreasing velocity.
@@
In Fig. 3, we show the position and velocity of the BH in our test setup for collisionless matter. As expected, the velocity of the BH decreases over the course of the simulation. The results for two different box sizes are shown. We find that the dynamical friction is reduced for a smaller box size. This is not surprising, as $r_{\mathrm{max}}$ in the Coulomb logarithm of the Chandrasekhar formula, Eq. (7), depends for our setup effectively on the box size because the box size limits the maximal possible impact parameter. The results agree qualitatively with the Chandrasekhar formula, after an initial phase during which the density wake forms, implying that Eq. (7) provides a good description when the Coulomb logarithm is taken as a free parameter. For the large box, we show the results using two different resolutions. In both cases, the trajectory of the BH is very small, such that we can safely assume that the resolution for our study is large enough.
Figure 4: As in Fig. 3, position and velocity of the BH as a function of time. The high-resolution setup with the large box size of $l=829.5\mu\mathrm{pc}$ is displayed. We fitted Chandrasekhar’s formula (Eq. 6 and 7) to the simulation results to determine the Coulomb logarithm, $\ln(\Lambda)$, with $\Lambda=r_{\mathrm{max}}/r_{\mathrm{min}}$. Besides the simulation result (dotted lines), we also show the fitted curves (solid lines).
We also quantitatively determined the Coulomb logarithm, $\ln(\Lambda)=\ln(r_{\mathrm{max}}/r_{\mathrm{min}})$, by fitting $\Lambda$ to the simulation data for the run with the high resolution ($N_{\mathrm{DM}}=5\times 10^{6}$) and the large box size ($l=829.5\mu\mathrm{pc}$). Here we did not use all of the simulation data but exclude the first phase of the simulation where the density wake forms and used the BH velocities for the fit only. The simulation results together with the fitted curves are displayed in Fig. 4. Our fit yielded a value of $\Lambda_{\mathrm{fit}}\approx 6.4$. Besides we can analytically estimate the expected value for $\Lambda$. It can be expressed as (Binney & Tremaine, 2008)
$\Lambda=\frac{b_{\mathrm{max}}\,v^{2}_{\mathrm{typ}}}{\mathrm{G}\,M}.$ (15)
We approximated the maximum impact parameter, $b_{\mathrm{max}}$, with half the size of the box length, $l$. The typical velocity, $v_{\mathrm{typ}}$, between the BH and the DM particles is approximately $\approx 10^{3}\,\mathrm{km}\,\mathrm{s}^{-1}$. The mass, $M$, of the perturber might effectively be larger than the BH mass as DM particles are bound to the BH. We discuss this in Sect. 4.3. Here we can assume, $M\approx 2\times m_{\mathrm{BH}}$. Last, $\mathrm{G}$, denotes the gravitational constant. Based on these assumptions we obtained a value of $\Lambda\approx 6.9$, which agrees well with our simulation result ($\Lambda_{\mathrm{fit}}\approx 6.4$), especially given that the logarithm of that value matters for the dynamical friction.